### v2

@@ Abstract
Turbulent-flow control aims to develop strategies that effectively manipulate fluid systems, such as the reduction of drag in transportation and enhancing energy efficiency, both critical steps towards reducing global CO2 emissions. Deep reinforcement learning (DRL) offers novel tools to discover flow-control strategies, which we combine with our knowledge of the physics of turbulence. We integrate explainable deep learning (XDL) to objectively identify the coherent structures containing the most informative regions in the flow, with a DRL model trained to reduce them. The trained model targets the most relevant regions in the flow to sustain turbulence and produces a drag reduction which is higher than that of a model specifically trained to reduce the drag, while using only half its power consumption. Moreover, the XDL model results in a better drag reduction than other models focusing on specific classically identified coherent structures. This demonstrates that combining DRL with XDL can produce causal control strategies that precisely target the most influential features of turbulence. By directly addressing the core mechanisms that sustain turbulence, our approach offers a powerful pathway towards its efficient control, which is a long-standing challenge in physics with profound implications for energy systems, climate modeling and aerodynamics.
@@
Turbulence remains one of the last unsolved problems of classical physics [1], defying a complete theoretical description despite its ubiquitous presence in natural [2] and engineering [3] flows. In fact, turbulent fluid flows are present in a myriad of industrial applications: from the food industry to the cooling of microprocessors, or harnessing wind power. Turbulence is also the primary cause of viscous drag with $30\%$ of the energy consumption worldwide being spent on overcoming drag in transportation [4]. Thus, improving the aerodynamic efficiency of terrestrial and airborne vehicles plays a pivotal role in the global reduction of CO2 emissions. Despite centuries of theoretical [5], experimental [6], and computational [7] progress since Leonardo da Vinci’s first observations [8], many fundamental questions on the nature of turbulence remain unsolved [9, 10, 11], preventing us from efficiently manipulating turbulent flows.
@@
Figure 1: The overall architecture of the deep reinforcement learning framework. The scheme shows the communication between the involved parts. The top shows the information inputs to the DRL agent, i.e. the rewards ($r_{t}$) and states ($s_{t}$) at each time $t$, and outputs to the flow, i.e. the actions ($a_{t}$). In this multi-agent reinforcement-learning (MARL) framework, the agents cooperate featuring a shared neural network. The middle panel depicts the computational fluid dynamics (CFD) environment, from which each agent receives the information $(u,v)$ of the point above within a plane located at $y^{+}=15$. The actions are applied at the wall in the form of blowing and suction. At the bottom, we illustrate the different rewards considered. We note that the SHAP values are computed instantaneously from the velocity field through a pre-trained U-net model.
@@
In this work, we incorporate the most recent progress in understanding the underlying dynamics of turbulence to design novel control strategies which reduce the friction produced by the flow. In the chaotic motion that characterizes wall-bounded turbulence, it is possible to identify persistent spatio-temporal patterns (coherent structures) that interact in the near-wall cycle of turbulence [12, 13]. This cycle involves streamwise vortices generating elongated velocity perturbations (streaks), which subsequently experience instabilities and breakdown, generating new vortices and giving rise to a self-sustaining process (SSP) [14]. Existing approaches to drag reduction, such as opposition control [15], seek to disrupt this SSP based on heuristic observations, for instance by targeting ejection and sweep events [16]. Recently, the advent of new data-intensive techniques has expanded the available toolbox, with deep reinforcement learning (DRL) emerging as a particularly powerful approach to reveal new, more effective control strategies. DRL has demonstrated success across diverse physical domains, as broad as plasma fusion [17], optics [18], and fluid dynamics. Within this domain, DRL has been successfully applied to a wide range of control problems, including coordinated motion in fish schooling [19] and the control of separated flows [20, 21].
@@
The framework for DRL control is illustrated in Fig. 1: the DRL agent (or controller) receives partial observations $s_{t}$ of the system and is trained to generate actions $a_{t}$, which modify the turbulent flow in the effort to maximize the reward $r_{t}$. While previous DRL applications to turbulence control have targeted directly the reduction of turbulent drag [22, 23], our approach targets the SSP of near-wall turbulence, with rewards based on the agent’s ability to manipulate turbulent coherent structures.
@@
We identify these structures using both conventional criteria (c.f. [16] and references therein) and data-driven methods based on explainable deep learning (XDL) [24]. Conventional coherent structures in turbulence, i.e. Q events, streaks, and vortices, are characterized using instantenous velocity fluctuations and their derivatives, while data-driven structures are identified using Shapley additive explanation (SHAP) values [25], which are obtained through a game-theoretic method that calculates the importance of each input feature in a deep neural network. When applied to a U-net based on convolutional layers used for prediction of the future states of the flow, the SHAP values successfully identify regions of importance in wall-bounded turbulence [24].
@@
The present study demonstrates how XDL can effectively be coupled with DRL to improve the performance of the learned policies compared with conventional physics-based rewards. We evaluate four reward strategies: (i) wall-shear stress reduction, (ii) Q-event reduction, (iii) streak reduction, and (iv) reduction of the SHAP values. The first reward is the most direct approach, rewards (ii) and (iii) target coherent structures traditionally studied in the turbulence literature, while (iv) is entirely data-driven yet physics-aware via XDL. To design the rewards, we choose appropriate instantaneous proxies to characterize the intensity of the turbulent coherent structures. In the SHAP-based reward, a U-net is used to predict the SHAP values from the instantaneous velocity field, as discussed in more detail in the Methods section.
@@
Our work presents an entirely data-driven method that requires no prior flow hypotheses, and yet is capable of identifying the physical processes governing wall-bounded turbulence dynamics. The implications of the present study extend beyond turbulence control, suggesting a powerful tool where XDL and DRL can uncover fundamental physical mechanisms that might otherwise remain out of reach using traditional approaches.
@@
Figure 2: Reduction of the quantities of interest with respect to the uncontrolled case. (Left) Drag reduction, (top-right) proxy for Q-event reduction and (bottom-right) proxy for streak reduction. Results are averaged over the 50 initial conditions used for policy evaluation. Solid lines denote mean values and shaded regions indicate one standard deviation. Colors indicate different control strategies: opposition control (blue), DRL for direct drag reduction (orange), DRL for Q-event reduction (green), DRL for streak reduction (red), and DRL for SHAP reduction (purple).
Our DRL set-up strategically identifies coherent structures to effectively influence the SSP of near-wall turbulence. This approach requires quantitative proxies to characterize the coherent structures in the flow. We define the instantaneous velocity vector $\mathbf{u}(x,y,z,t)=(u,v,w)$ along the streamwise, wall-normal and spanwise directions, where $t$ denotes time. Deviations from the instantaneous spatial mean averaged in the periodic directions are represented by primed variables: $u^{\prime}(x,y,z,t)=u(x,y,z,t)-\iint u(x,y,z,t)\mathrm{d}x\mathrm{d}z/(L_{x}L_{z})$ and $L_{x}$ and $L_{z}$ denote the domain sizes. We consider a turbulent open channel which can be characterized by the friction Reynolds number $Re_{\tau}=u_{\tau}h/\nu$, where $\nu$ denotes the fluid kinematic viscosity, $h$ the channel height, and $u_{\tau}=\sqrt{\tau_{w}/\rho}$ is the friction velocity (defined in terms of the wall-shear stress $\tau_{w}$ and the fluid density $\rho$). We consider $Re_{\tau}=180$ in our uncontrolled simulations. Quantities non-dimensionalized with the viscous scales $u_{\tau}$ and $\nu$ are denoted by the superscript ``+"``"``+"` ` + ". The DRL-discovered control strategies will be compared with the heuristic opposition control. Opposition control is a closed-loop control, where the control action is based on real-time flow-state sensing [15], which reduces friction drag by using blowing and suction with a vertical velocity distribution defined as:
@@
This control opposes the vertical velocity on a plane at a wall-normal distance $y_{s}$, with the aim of suppressing the streamwise vortices.
The following control inputs are sampled at $y^{+}=15$: $v^{\prime}(t,x,y^{+}=15,z)$ for opposition control and $\{u^{\prime}(t,x,y^{+}=15,z),v^{\prime}(t,x,y^{+}=15,z)\}$ for the DRL-based techniques. To quantify the total intensity of Q events we consider the Reynolds stress averaged over the domain volume:
@@
where $U_{T}$ is the long-time spatio-temporal average for the turbulent uncontrolled case and $\Omega$ denotes the volume. Following [26], we characterize the streak intensity as:
@@
where $\bm{\phi}=(\phi_{x},~{}\phi_{y},~{}\phi_{z})$ represent the SHAP vector field as detailed in the Methods section. Here, the SHAP values are inferred instantaneously from the velocity field by a U-net trained on the uncontrolled channel data (see the Methods section for more details).
@@
We consider two channel domain sizes in our study, one for training and one for testing. The DRL agents are trained in a small channel configuration (SCC), the so-called minimal flow unit [27], which is computationally affordable and supports a single SSP. The learnt policy is then tested in a large channel configuration (LCC) where many SSPs are present and interact. Here we discuss the performance in the larger channel, while the details of the computational settings and the SCC results are reported in the Methods section and in the Supplementary material, respectively.
@@
The discovered policies exhibit significant differences between the various proposed rewards. Fig. 2 shows various performance metrics averaged over 50 different initial conditions. Our results exhibit a significant drag reduction: approximately $31.7\%$ when using DRL aiming specifically at direct drag (DD) reduction, $27.5\%$ when targeting Q events, and $21.2\%$ when targeting streaks. However, all these approaches show a worse performance than the one of the SHAP-based policy, which presents a $33.3\%$ drag reduction: $5\%$ more (1.6 percentage points) than DD-based control, $20.8\%$ more (5.6 percentage points) than Q-event-based control, and $56.8\%$ more (12.1 percentage points) than streak-based control. This finding is particularly remarkable given that all policies were initially trained in a small channel where performance differences were small (see the Methods section and Supplementary material).
@@
We obtain further insights into the dynamic effect of the DRL-discovered control strategies when comparing the averaged Reynolds stresses and streak intensity in the various controlled cases, as shown in Fig. 2 (right). Note that the DRL approach targeting Q events exhibits the best performance reducing both Q events and streaks. Interestingly, the SHAP-based policy yields the highest drag reduction, despite not being particularly effective at reducing Q events or streaks, highlighting the fact that these classically studied structures are not the most important to target when reducing drag. The effectiveness of the SHAP-based policy stems from its ability to identify the most informative flow regions to predict future turbulence fluctuations in an objective manner. The DRL approach targeting SHAP values manipulates the most relevant structures for the flow evolution and the SSP of near-wall turbulence. By targeting these regions, the DRL control learns to hamper turbulence-supporting mechanisms in a causally informed manner, unlike the more simplistic approach of targeting the classically-studied coherent structures. These findings confirm that the traditional coherent structures provide an incomplete picture of the mechanisms supporting turbulence, and show that XDL complements and expands our understanding of wall-bounded turbulence [28].
@@
Figure 3: Joint probability density function of the streamwise and wall-normal fluctuations at $y^{+}=15$ for the different control strategies. Results are averaged over the 50 initial conditions used for policy evaluation for: uncontrolled flow (black), opposition control (blue), DRL for DD (orange), DRL for Q events (green), DRL for streaks (red), and DRL for SHAP values (purple). Grey lines in all the panels show the uncontrolled case for comparison.
@@
Fig. 3 shows the joint probability density function of streamwise ($u^{\prime}$) and wall-normal ($v^{\prime}$) perturbations on the sensing plane $y^{+}=15$ evaluated on 50 different trajectories. Note that the predominance of ejections and sweeps in the uncontrolled case is consistent with the wall-bounded turbulence literature [29] and the results from opposition control and DRL for drag reduction agree with our previous results [22]. Furthermore, the approaches based on classical structures produce distinctly different fluctuations intensities: Q-event-based control substantially reduces the perturbation intensities in both stream- and spanwise directions, while the streak-based control decreases streamwise perturbations but increases wall-normal perturbations, emphasizing the negative fluctuations. In contrast, our SHAP-based control leads to a quite even distribution of the four quadrants, with a slight preference towards positive streamwise fluctuations.
@@
Figure 4: Power input and time signal of a single controller for the various control strategies. (Top left) Power input averaged over the 50 initial conditions used for policy evaluation. (Rest of panels) Sample time signals for a single actuator in: opposition control (blue), DRL for drag reduction (orange), DRL for Q events reduction (green), DRL for streak reduction (red), and DRL for SHAP reduction (purple).
@@
An additional advantage of our XDL approach is observed when analyzing control dynamics and energy requirements. The power input of the control is defined as:
$w_{\text{in}}=\frac{1}{2}||v_{\text{wall}}||^{3}$ (5)
@@
for each agent. Fig. 4 shows sample actuation signals from the different control strategies, along with their associated power input requirements. Direct drag reduction control and streak-based control frequently oscillate between the minimum and maximum allowed control values in a “bang-bang” pattern, resulting in high power input. The SHAP-based control operates more efficiently, rarely reaching maximum or minimum control values and exhibiting fewer extreme actuation patterns. This translates to considerable input power savings: the SHAP-based approach requires only half the power input compared with the DD control while delivering better drag reduction. The Q-event-based control, while even more energy-efficient, yields significantly lower drag reduction than the SHAP-based approach, as discussed above.
@@
$S=\frac{c_{f,\text{uncontrolled}}-(c_{f}+w_{\text{in}})}{c_{f,\text{uncontrolled}}},$ (6)

### v3

@@ Abstract
Turbulent-flow control aims to develop strategies that effectively manipulate fluid systems, such as the reduction of drag in transportation and enhancing energy efficiency, both critical steps towards reducing global CO2 emissions. Deep reinforcement learning (DRL) offers novel tools to discover flow-control strategies, which we combine with our knowledge of the physics of turbulence. We integrate explainable deep learning (XDL) to objectively identify the coherent structures containing the most informative regions in the flow, with a DRL model trained to reduce them. The model trained with XDL targets the most relevant regions in the flow to sustain turbulence and produces a drag reduction which is higher than that of a model specifically trained to reduce the drag, resulting in a $18.1\%$ better net-energy saving. The XDL-based control remains the most effective control strategy when generalizing across Reynolds numbers and geometries. This demonstrates that combining DRL with XDL can produce causal control strategies that precisely target the most influential features of turbulence. By directly addressing the core mechanisms that sustain turbulence, our approach offers a powerful pathway towards its efficient control, which is a long-standing challenge in physics with profound implications for energy systems, climate modeling and aerodynamics.
@@
Turbulence remains one of the last unsolved problems of classical physics [frisch1995turbulence], defying a complete theoretical description despite its ubiquitous presence in natural [ramalingam1960convective] and engineering [spillane1967clear] flows. In fact, turbulent fluid flows are present in a myriad of industrial applications: from the food industry to the cooling of microprocessors, or harnessing wind power. Turbulence is also the primary cause of viscous drag with $30\%$ of the energy consumption worldwide being spent on overcoming drag in transportation [IEA2020]. Thus, improving the aerodynamic efficiency of terrestrial and airborne vehicles plays a pivotal role in the global reduction of CO2 emissions and would have enormous economical consequences, since a reduction in drag achieving a fuel reduction on truck fleets of $5\%$ would correspond to an increase in profit of over $100\%$[brunton2015closed]. Despite centuries of theoretical [marusic_science], experimental [avila_science], and computational [cardesa_science] progress since Leonardo da Vinci’s first observations [marusic2021leonardo], many fundamental questions on the nature of turbulence remain unsolved [gibson1962spectra, argoul1989wavelet, greenstein1970superfluid], preventing us from efficiently manipulating turbulent flows.
@@
Figure 1: The overall architecture of the deep reinforcement learning framework. The scheme shows the communication between the involved parts. The top shows the information inputs to the DRL agent, i.e. the reward ($r_{t}$) and state ($s_{t}$) at each time $t$, and outputs to the flow, i.e. the action ($a_{t}$). In this multi-agent reinforcement-learning (MARL) framework, the agents cooperate featuring a shared neural network. The middle panel depicts the computational fluid dynamics (CFD) environment, from which each agent receives the information $(u,v)$ of the point above within a plane located at $y^{+}=15$. The actions are applied at the wall in the form of blowing and suction. At the bottom, we illustrate the different rewards considered. We note that the SHAP values are computed instantaneously from the velocity field through a pre-trained U-net model.
@@
In this work, we incorporate the most recent progress in understanding the underlying dynamics of turbulence to design novel control strategies which reduce the friction produced by the flow. In the chaotic motion that characterizes wall-bounded turbulence, it is possible to identify persistent spatio-temporal patterns (coherent structures) that interact in the near-wall cycle of turbulence [jimenez1999autonomous, jimenez2013near]. This cycle involves streamwise vortices generating elongated velocity perturbations (streaks), which subsequently experience instabilities and breakdown, generating new vortices and giving rise to a self-sustaining process (SSP) [waleffe1997self]. Existing approaches to drag reduction, such as opposition control [choi1994active], seek to disrupt this SSP based on heuristic observations, for instance by targeting so-called ejection and sweep events [jimenez2018coherent]. Recently, the advent of new data-intensive techniques has expanded the available toolbox, with deep reinforcement learning (DRL) emerging as a particularly powerful approach to reveal new, more effective control strategies. DRL has demonstrated success across diverse physical domains, as broad as plasma fusion [seo2024avoiding], optics [nousiainen2021adaptive], and fluid dynamics. Within this domain, DRL has been successfully applied to a wide range of control problems, including coordinated motion in fish schooling [gazzola2014reinforcement] and the control of separated flows [gautier2015closed, font2025deep].
@@
The framework for DRL control is illustrated in Fig. 1: the DRL agent (or controller) receives partial observations $s_{t}$ of the system and is trained to generate actions $a_{t}$, which modify the turbulent flow in the effort to maximize the reward $r_{t}$. While previous DRL applications to turbulence control have targeted directly the reduction of turbulent drag [guastoni2023deep, Sonoda_Liu_Itoh_Hasegawa_2023], our approach targets the SSP of near-wall turbulence, with rewards based on the agent’s ability to manipulate turbulent coherent structures.
@@
We identify these structures using both conventional criteria (c.f. [jimenez2018coherent] and references therein) and data-driven methods based on explainable deep learning (XDL) [cremades2024identifying]. Conventional coherent structures in turbulence, i.e. Q events, streaks, and vortices, are characterized using instantenous velocity fluctuations and their derivatives, while data-driven structures are identified using Shapley additive explanation (SHAP) values [lundberg2017unified], which are obtained through a game-theoretic method that calculates the importance of each input feature in a deep neural network. This idea has been successfully applied to explain relaminarization events in wall-bounded shear flows [lellep2022interpreted], where a XGBoost model trained to predict relaminarization events performed best when trained on features identified by the SHAP values. When applied to a U-net based on convolutional layers used for prediction of the future states of the flow, the SHAP values have successfully identified regions of importance in wall-bounded turbulence [cremades2024identifying].
@@
The present study demonstrates how XDL can effectively be coupled with DRL to improve the performance of the learned policies compared with conventional physics-based rewards. We will firstly show the strength of the XDL and DRL combination on a two-dimensional illustrative problem, moving afterwards towards numerical simulations of turbulent flows. In the flow problem we will evaluate four reward strategies: (i) wall-shear stress reduction, (ii) Q-event reduction, (iii) streak reduction, and (iv) reduction of the SHAP values. The first reward is the most direct approach, rewards (ii) and (iii) target coherent structures traditionally studied in the turbulence literature, while (iv) is entirely data-driven yet physics-aware via XDL. To design the rewards, we choose appropriate instantaneous proxies to characterize the intensity of the turbulent coherent structures. In the SHAP-based reward, a U-net is used to predict the SHAP values from the instantaneous velocity field, as discussed in more detail in the Methods section.
@@
Our work presents an entirely data-driven method that is free from assumptions about which coherent structures are most relevant for turbulence, and yet is capable of identifying the physical processes governing wall-bounded turbulence dynamics. Should we consider an entirely different flow problem, the SHAP values would need to be recomputed. However, in the present work we show that predicting them through a U-net trained to predict SHAP values from an uncontrolled velocity field yields good results for control purposes. The implications of the present study extend beyond turbulence control, suggesting a powerful tool where XDL and DRL can uncover fundamental physical mechanisms that might otherwise remain out of reach using traditional approaches.
@@
Figure 2: Reduction of the quantities of interest with respect to the uncontrolled case. (Left) Drag reduction, (top-right) proxy for Q-event reduction and (bottom-right) proxy for streak reduction. Results are averaged over the 150 initial conditions used for policy evaluation. Solid lines denote mean values and shaded regions indicate one standard deviation. Colors indicate different control strategies: opposition control (blue), DRL for direct drag reduction (orange), DRL for Q-event reduction (green), DRL for streak reduction (red), and DRL for SHAP reduction (purple).
Our DRL set-up strategically identifies coherent structures to effectively influence the SSP of near-wall turbulence. This approach requires quantitative proxies to characterize the coherent structures in the flow. We define the instantaneous velocity vector $\mathbf{u}(x,y,z,t)=(u,v,w)$ along the streamwise, wall-normal and spanwise directions, where $t$ denotes time. Deviations from the instantaneous spatial mean averaged in the periodic directions are represented by primed variables: $u^{\prime}(x,y,z,t)=u(x,y,z,t)-\iint u(x,y,z,t)\mathrm{d}x\mathrm{d}z/(L_{x}L_{z})$ and $L_{x}$ and $L_{z}$ denote the domain sizes. We consider a turbulent open channel which can be characterized by the friction Reynolds number $Re_{\tau}=u_{\tau}h/\nu$, where $\nu$ denotes the fluid kinematic viscosity, $h$ the channel height, and $u_{\tau}=\sqrt{\tau_{w}/\rho}$ is the friction velocity (defined in terms of the wall-shear stress $\tau_{w}$ and the fluid density $\rho$). We consider $Re_{\tau}=180$ in our uncontrolled simulations. Quantities non-dimensionalized with the viscous scales $u_{\tau}$ and $\nu$ are denoted by the superscript +``+". The DRL-discovered control strategies will be compared with the heuristic opposition control. Opposition control is a closed-loop control, where the control action is based on real-time flow-state sensing [choi1994active], which reduces friction drag by using blowing and suction at the wall with a vertical velocity distribution defined as:
@@
This control opposes the vertical velocity on a plane at a wall-normal distance $y_{s}$, with the aim of suppressing the streamwise vortices. The following control inputs are sampled at $y^{+}=15$: $v^{\prime}(t,x,y^{+}=15,z)$ for opposition control and $\{u^{\prime}(t,x,y^{+}=15,z),v^{\prime}(t,x,y^{+}=15,z)\}$ for the DRL-based techniques. To quantify the total intensity of Q events we consider the Reynolds stress averaged over the domain volume:
@@
where $U_{T}$ is the long-time spatio-temporal average for the turbulent uncontrolled case and $\Omega$ denotes the volume. Following [kline1967structure], we characterize the streak intensity as:
@@
where $\bm{\phi}=(\phi_{x},~\phi_{y},~\phi_{z})$ represent the SHAP vector field as detailed in the Methods section. Here, the SHAP values are inferred instantaneously from the velocity field by a U-net trained on the uncontrolled channel data (see the Methods section for more details). Note that these proxies only quantify the magnitude of the relevant coherent structure. We train our DRL agents to reduce these coherent structures without considering a priori wether they might be beneficial or not for drag reduction.
@@
We consider two channel domain sizes in our study, one for training and one for testing. The DRL agents are trained in a small channel configuration (SCC), the so-called minimal flow unit [jimenez1991minimal], which is computationally affordable and supports a single SSP. The learnt policy is then tested in a large channel configuration (LCC) where many SSPs are present and interact. The number of agents is always kept equal to the number of gridpoints at the wall in our simulations, i.e. the evaluation of the policy in the LCC is achieved by replicating the learnt control. Here we discuss the performance in the larger channel, while the details of the computational settings and the SCC results are reported in the Methods section and in the Supplementary material, respectively.
@@
The discovered policies exhibit significant differences between the various proposed rewards. Fig. 2 shows various performance metrics averaged over 150 different initial conditions. Our results exhibit a significant drag reduction: approximately $31.9\%$ when using DRL aiming specifically at direct drag (DD) reduction, $26.9\%$ when targeting Q events, and $20.9\%$ when targeting streaks. However, all these approaches show a worse performance than that of the SHAP-based policy, which presents a $33.7\%$ drag reduction: $5.6\%$ more (1.8 percentage points) than DD-based control, $25.3\%$ more (6.8 percentage points) than Q-event-based control, and $61.2\%$ more (12.8 percentage points) than streak-based control. This finding is particularly remarkable given that all policies were initially trained in a small channel where performance differences were small (see the Methods section and Supplementary material). We note the strong periodicity of the time signal, which could be smoothed by considering a shorter actuation $\Delta t$ (not shown). This is further illustrated in the zero-shot generalization for $Re_{\tau}=550$, where the actuation frequency remains the same in inner units, but is increased in outer units. It might appear counter-intuitive that a model trained to reduce the SHAP values outperforms a model trained for direct drag reduction when reducing the drag. However, the system under consideration is not entirely Markovian due to partial observability and the SHAP values can be seen as sophisticated reward shaping. A more detailed discussion using an illustrative navigation problem can be found in the Supplementary material.
@@
We obtain further insights into the dynamic effect of the DRL-discovered control strategies when comparing the averaged Reynolds stresses and streak intensity in the various controlled cases, as shown in Fig. 2 (right). Note that the DRL approach targeting Q events exhibits the best performance reducing both Q events and streaks. Interestingly, the SHAP-based policy yields the highest drag reduction, despite not being particularly effective at reducing Q events or streaks, highlighting the fact that these classically studied structures are not the most important to target when reducing drag. The effectiveness of the SHAP-based policy stems from its ability to identify the most informative flow regions to predict future turbulence fluctuations in an objective manner. The DRL approach targeting SHAP values manipulates the most relevant structures for the flow evolution and the SSP of near-wall turbulence. By targeting these regions, the DRL control learns to hamper turbulence-supporting mechanisms in a causally informed manner, unlike the more simplistic approach of targeting the classically-studied coherent structures. These findings confirm that the traditional coherent structures provide an incomplete picture of the mechanisms supporting turbulence, and show that XDL complements and expands our understanding of wall-bounded turbulence [cremades2024classically].
@@
Figure 3: Joint probability density function of the streamwise and wall-normal fluctuations at $y^{+}=15$ for the different control strategies. Results are averaged over the 150 initial conditions used for policy evaluation for: uncontrolled flow (black), opposition control (blue), DRL for $D$ (orange), DRL for Q events (green), DRL for streaks (red), and DRL for SHAP values (purple). Grey lines in all the panels show the uncontrolled case for comparison.
@@
Fig. 3 shows the joint probability density function of streamwise ($u^{\prime}$) and wall-normal ($v^{\prime}$) perturbations on the sensing plane $y^{+}=15$ evaluated on 150 different trajectories. Note that the predominance of ejections and sweeps in the uncontrolled case is consistent with the wall-bounded turbulence literature [lozano2012three] and the results from opposition control and DRL for drag reduction agree with our previous results [guastoni2023deep]. Furthermore, the approaches based on classical structures produce distinctly different fluctuations intensities: Q-event-based control substantially reduces the perturbation intensities in both stream- and wall-normal directions, while the streak-based control decreases streamwise perturbations but increases wall-normal perturbations, emphasizing the negative values of the fluctuations. In contrast, our SHAP-based control leads to a quite even distribution of the four quadrants, with a slight preference towards positive streamwise fluctuations. Note that we are exclusively reporting the effect of the control on the velocity perturbations of the flow. The effect of each particular control on the drag reduction is highlighted in Fig. 2. Additional insights into the effects of the various control strategies are included in the Supplementary material, which includes flow visualizations and spectral analyses.
@@
Figure 4: Power input and time signal of a single controller for the various control strategies. (Top left) Total power input averaged over the 150 initial conditions used for policy evaluation. (Rest of panels) Sample time signals for a single actuator in: opposition control (blue), DRL for drag reduction (orange), DRL for Q events reduction (green), DRL for streak reduction (red), and DRL for SHAP reduction (purple). The actuator location is the same in all the cases: $(x_{0},y_{0},z_{0})=(0,0,0)$.
@@
An additional advantage of our XDL approach is observed when analyzing control dynamics and energy requirements. The total power input of the control is defined as:
[…]
@@
for each agent, where the first term denotes the kinematic power input, and $S_{1,2}=0$ if $v\leq 0$ (or $pv\leq 0$), and $S_{1,2}=1$ if $pv>0$ following [hasegawa2011dissimilar, Sonoda_Liu_Itoh_Hasegawa_2023] to ensure that no energy recovery from the pressure is considered. Fig. 4 shows sample actuation signals from the different control strategies, along with their associated power input requirements. Direct drag reduction control frequently oscillates between the minimum and maximum allowed control values in a “bang-bang” pattern, resulting in high power input. The SHAP-based control and streak-based control operate more efficiently, rarely reaching maximum or minimum control values and exhibiting fewer extreme actuation patterns. This translates to considerable input power savings: the SHAP-based approach requires only half the power input compared with the DD control while delivering better drag reduction. The Q-event-based control, while even more energy-efficient, yields significantly lower drag reduction than the SHAP-based approach, as discussed above.
@@
$S=\frac{c_{f,\text{uncontrolled}}-(c_{f}+W)}{c_{f,\text{uncontrolled}}},$ (6)
@@
Figure 5: Illustration of the Z zero-shot generalization process. (Top left) Illustration of the learnt policies and training in SCC at $\textit{Re}_{\tau}=180$, red and blue contours indicate illustrative values of $v$. (Top right) Evaluation of the policy on a large channel (LCC) at $\textit{Re}_{\tau}=180$, red and blue contours indicate illustrative values of $v$. (Bottom) Results from the zero-shot application of the policies learnt on the SCC at $\textit{Re}_{\tau}=180$ on a turbulent channel flow at $\textit{Re}_{\tau}=550$ and a turbulent boundary layer with a $\textit{Re}_{\theta}=460$, red and blue contours indicate illustrative values of $w$. These flow cases are detailed in the Methods section and 3D visualizations for the uncontrolled cases are added to illustrate the flow complexity. In all cases the SHAP-based policy is the best performing one.
Zero-shot generalization
Zero-shot deployment of a DRL has proved to be successful even in complex geometries, such as the flow around wings [wang2026physics]. We further assess the drag-reduction strategies discovered in this work through a zero-shot generalization to two other flow cases: (i) a channel flow at $\textit{Re}_{\tau}=550$ and (ii) zero-pressure-gradient turbulent boundary layer (ZPGTBL).
Firstly, we consider the turbulent flow in the LCC at a $\textit{Re}_{\tau}=550$ instead of the training one of $\textit{Re}_{\tau}=180$. Our DRL policy is deployed in this configuration by keeping the same set-up in inner units, i.e. sensing plane remains at the same $y^{+}=15$, that the amplitude of the control remains $u_{\tau}$, and the control frequency remains at the same $\Delta t^{+}\approx 0.54$, for consistency, we keep the observations normalized with the original $u_{\tau}$ from training. The results for the drag reduction averaged over 150 different initial conditions corresponding to the various DRL policies shown in Fig. 5. We observe that while all policies degrade with respect to the $\textit{Re}_{\tau}=180$ case, the SHAP-based policy still outperforms all other policies resulting in a drag reduction of $23.5\%$, which is $4.3\%$ better (1 percentage point) than the DD-based control, $29.8\%$ better (5.4 percentage points) than the Q events-based control, $42.4\%$ better (7 percentage points) than the streaks-based control and $77.4\%$ better (10.25 percentage points) than opposition control. This shows that the SHAP-based control effectively generalises well beyond its training configuration (as long as the SHAP structures observed in the flow remain similar).
Secondly, we consider a ZPGTBL with $\textit{Re}_{\delta_{0}^{*}}=450$ at the inflow. We apply the control between $x\in[200,300]$ downstream of the inflow, where the turbulence fully developed [schlatter2009turbulent, guastoni2025fully] (more details can be found in the Methods section). In this case, since the flow is spatially developing, we use the uncontrolled streamwise-averaged quantities in inner units within the control region for our setup. Fig. 5 shows the spatial distribution of the $u_{\tau}$ along the boundary layer averaged in time, in the spanwise direction, and over 150 different initial conditions. The results show that in this configuration the DRL trained to minimize the SHAP values in the flow yields the lowest $u_{\tau}$. The drag reduction obtained within the control region is $32\%$ for the SHAP-based control, this is $295.1\%$ better (23.9 percentage points) than the DD-based control, $316\%$ better (24.3 percentage points) than the Q events-based control and $107.8\%$ (16.6 percentage points) better than opposition control. We note that the streaks-based control only produces $0.05\%$ drag reduction in this case, being mostly inefficient. It is worth noting that opposition control is better than all of the DRL-discovered control strategies, except for the SHAP-based control.
These zero-shot generalizations are applied to flow configurations significantly different from the training set-up to illustrate the flexibility and robustness achieved through XDL for turbulence control.